OpenStudy (anonymous):

Factor completely: 18x^2 − 21x − 15

OpenStudy (nathan917):

Factor the expression by the GCF of 3: 3(6x² - 7x - 5) Then, by factoring decomposition: 3(6x² + 3x - 10x - 5) = 3(3x(2x + 1) - 5(2x + 1)) = 3(2x + 1)(3x - 5) @AriPotta this better?

OpenStudy (aripotta):

well i don't think she knows what "factoring decomposition" is

OpenStudy (anonymous):

3x^2 + 12x + 7

OpenStudy (anonymous):

@nathan917

OpenStudy (aripotta):

if she truly understood how to do these problems, she wouldn't be posting them

OpenStudy (aripotta):

she'd be doing them herself

OpenStudy (nathan917):

-3x^2 + 12x = -3x(x - 4) = 3x(4 - x) i gtg bye!